# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

### v1

@@ Abstract
Novel metaphor comprehension involves complex semantic processes and linguistic creativity, making it an interesting task for studying language models (LMs). This study investigates whether surprisal, a probabilistic measure of predictability in LMs, correlates with different metaphor novelty datasets. We analyse surprisal from 16 LM variants on corpus-based and synthetic metaphor novelty datasets. We explore a cloze-style surprisal method that conditions on full-sentence context. Results show that LMs yield significant moderate correlations with scores/labels of metaphor novelty. We further identify divergent scaling patterns: on corpus-based data, correlation strength decreases with model size (inverse scaling effect), whereas on synthetic data it increases (Quality–Power Hypothesis). We conclude that while surprisal can partially account for annotations of metaphor novelty, it remains a limited metric of linguistic creativity.
@@
Omar Momen,   Emilie Sitter, Berenike Herrmann, and Sina Zarrieß CRC 1646 – Linguistic Creativity in Communication Faculty of Linguistics and Literary Studies Bielefeld University, Germany {omar.hassan,emilie.sitter,berenike.herrmann,sina.zarriess}@uni-bielefeld.de
@@ 1 Introduction
Recent advances in neural language modelling have led to a renewed interest in studying linguistic creativity, an aspect of language that was notoriously challenging for traditional NLP systems. A well-known instance of creative language is metaphor, which arises through mapping of a source domain–a lexical unit’s literal meaning–onto a target domain–its figurative meaning-cmt1980. Yet, not all metaphors are equally novel or creative as many of such mappings are highly conventionalized, such as in the famous example “to attack an argument,” where the mapping from the source domain “WAR” onto the target domain “ARGUMENT” corresponds to a conventional sense of the verb “attack” which can be found in dictionaries. Other metaphors, as in “The arrested water” (Figure 1) establish a more unconventional or novel mapping that is considered creative.
@@
Figure 1: A sentence from VUA-ratings with metaphor-novelty annotations (n) and surprisal measures (s) from GPT2-base. The top part shows the direct-surprisal measure, while the bottom part shows the cloze-surprisal measure.
@@
The distinction between conventional and novel metaphors is well established in theory BowdleGentner2005; Gibbs_2006 and studied in experimental work on cognitive processing ArzouanGoldsteinFaust2007; lai2009. Novel metaphors require greater interpretative effort compared to conventional metaphors, as the unfamiliar mapping requires speakers to construct new connections between domains philip_2016_conventional_novel.
@@
However, separating novel metaphors from conventional ones remains a challenging task for human-based annotations do-dinh-etal-2018-weeding; reimann-scheffler-2024-metaphor. While computational methods on metaphor detection and understanding do not even take this distinction into account in most cases (maudslay-teufel-2022-metaphorical), and some existing studies highlight that novel metaphors are more difficult to detect than conventional ones neidlein-etal-2020-analysis; tong-etal-2024-metaphor. In this work, we propose to fill this gap and study questions on LMs processing of novel and conventional metaphors, and investigate LM-based metrics correlation to different setups of metaphor novelty annotations.
@@
We draw on a line of research that has studied the ability of LMs to account for effects of difficulty in human sentence processing and that goes back to surprisal theory (hale-2001-probabilistic; levy2008expectation). Surprisal is computed with LMs as the negative log probability of a word in context and has been found to provide a robust predictor for human processing difficulty (e.g., of reading times) (goodkind-bicknell-2018-predictive; shain2024large).
@@
However, recent work draws a mixed picture in terms of which LMs can provide the most robust and cognitively plausible predictors. oh-schuler-2023-surprisal observe an inverse scaling effect when testing surprisal estimates from GPT-2 models of different sizes, showing that the surprisal computed with smaller model sizes achieved a better fit with human reading times than larger model sizes. wilcox-etal-2023-language, on the other hand, train LMs of small and medium size on a range of languages and find that LM quality generally correlates with its psychometric predictive power.
@@
In this paper, we investigate surprisal on metaphor words computed with LMs as a metric of metaphor novelty. We find significant moderate correlations between LMs’ surprisal and different metaphor novelty annotations. We perform our experiments on four different metaphor novelty datasets coming from corpus-based and synthetic setups, using LMs of different sizes and architectures. Most interestingly, we observe effects supporting the "inverse scale" pattern oh-etal-2025-inverse on the two corpus-based novelty datasets, and contrary patterns supporting the counter-argument, "Quality-Power" hypothesis wilcox-etal-2023-language on the two synthetic datasets. We also investigate the effect of instruction-tuning on surprisal correlation to metaphor novelty. Moreover, we perform a deeper analysis on the level of genre splits, revealing that genre (the density of metaphors & LMs’ perplexity on that genre) can affect the quality of surprisal as a predictor of metaphor novelty significantly. Finally, we introduce a new variant of computing surprisal, cloze-surprisal, for recording word-level surprisal from causal LMs in a way that includes the right context of the word in its conditional probability. We find that this method can boost the correlation of surprisal to metaphor novelty annotations by a few points. In general, our study establishes a promising direction for studying linguistic creativity with LMs and calls for novel measures and datasets that provide systematic annotations of metaphor novelty across genres.
@@ 2.1 Metaphor Annotations
The most common annotation scheme in corpus-based metaphor studies is the Metaphor Identification Procedure Vrije Universiteit (MIPVU) mip2007; vuamc, designed as a reliable step-by-step framework for identifying metaphorically used words. This method was used to construct the VU Amsterdam Metaphor Corpus (VUAMC) vuamc, a large-scale, genre-balanced corpus containing 186,673 words sampled from fiction, news, academic writing, and conversations splits in the BNC Baby edition. The VUAMC has become a benchmark for metaphor research, inspiring developments in further English corpora beigman-klebanov-etal-2018-corpus; mohammad-etal-2016-metaphor and more recent multilingual efforts sanchez-bayona-agerri-2022-leveraging; egg-kordoni-2022-metaphor.
Several studies have extended annotation beyond binary literal/metaphoric labels. mohler-etal-2016-introducing introduced a four-point metaphoricity scale, considering factors such as vividness and familiarity. Reijnierse2018 proposed the concept of deliberateness, distinguishing metaphors intended to be recognized as metaphors; novel metaphors are typically deliberate under this framework. Direct novelty annotations were first introduced by parde_corpus_2018, who asked annotators to rate metaphorical word pairs from the VUAMC on a 0–3 novelty scale. do-dinh-etal-2018-weeding proposed a more comprehensive approach considering all metaphor words in VUAMC, and by aggregating ranked novelty judgments into continuous scores. More recently, reimann-scheffler-2024-metaphor proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries.
Psycholinguistic studies also produce datasets of metaphors, but these are smaller in size and feature controlled synthetic sentences (fixed words, structures, sentence length, etc.), cf. cardillo2010neural; cardillo2012selective; roncero2014role. Some of these also created datasets of metaphors that are classified as conventional or novel metaphors lai2009; ahrens-burgers-zhong-2024.
@@ 2.2 Predictive Powers of Surprisal
Surprisal estimates from LMs have been shown to capture human processing difficulty across multiple behavioural and neural measures. In self-paced reading and eye-tracking studies, token-level surprisals from causal models significantly predict reading times beyond lexical and syntactic factors goodkind-bicknell-2018-predictive; wilcox2020; oh-schuler-2023-transformer. Similar effects have been reported for acceptability judgments, where LM surprisals align with human sensitivity to gradient grammaticality lau-etal-2017; meister-etal-2021-revisiting; hu-etal-2020-systematic; tjuatja-etal-2025-goes.
@@ 2.3 LM-Based Metrics for Metaphor Novelty
Recent work has explored diverse LM-based approaches—prompting, surprisal, embedding similarity and attention patterns—for the analysis of metaphorical language. aghazadeh-etal-2022-metaphors demonstrate that metaphor-relevant information is encoded in mid-layer embeddings of multilingual pre-trained LMs. ichien2024largelanguagemodeldisplays show that GPT-4 generates interpretations of novel literary metaphors that are favoured by human judges over interpretations by college students, suggesting sensitivity to metaphorical meaning beyond lexical overlap. pedinotti-etal-2021-howling tested BERT’s masked token probabilities across conventional, novel, and nonsensical metaphors, showing that novel metaphors tend to receive lower probabilities, but distinctions between novelty and deviancy remained unclear. We extend this study towards more recent LMs and broader comparisons between models, datasets, and genres. djokic2021episodic trained a BERT-based classifier to predict novelty scores jointly with the task of metaphor detection.
@@ 3.1 Corpus‐based Datasets
do-dinh-etal-2018-weeding collected crowd-sourced ratings of metaphor novelty for VUAMC. In their set-up, annotators were presented four sentences containing an MRW (content words only) and asked to select the best (most novel) and worst (most conventional). Each MRW appeared in six different best-worst scaling comparisons. These annotations were then converted into continuous Best-Worst Scaling scores kiritchenko-mohammad-2017-best in the range of (-1, +1), with -1 being the most conventional and +1 being the most novel. Additionally, they also convert these scores to binary labels using a threshold of 0.5. This results in labelling 353 metaphors as novel out of 15,155 content metaphoric words in VUA.
@@
reimann-scheffler-2024-metaphor proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries. They applied this method to VUAMC. In particular, they re-annotated the potentially novel metaphors in VUA according to VUA-ratings and Reijnierse2018’s metaphor deliberateness annotations (1,160 potentially novel metaphors in total). Their procedure resulted in labelling only 409 content [^1] We exclude non-content words from this dataset to allow comparison with VUA-ratings, which originally excluded non-content words. metaphoric words as novel out of the 1,160 potentially novel metaphors. We assume the remaining metaphors in VUA to be conventional in our study.
@@ 3.2 Synthetic Datasets
lai2009 investigated how our brains handle conventional and novel metaphors differently. To this end, they constructed a set of controlled experimental items featuring metaphors with two degrees of novelty. Two linguists selected 104 words, and constructed 4 sentences for each word, according to the Conceptual Metaphor Theory (CMT) cmt1980. For each word, the items include one (i) literal use, (i) conventional metaphor, (iii) novel metaphor and (iv) anomalous use, with the target word as the last word in each sentence. Familiarity and interpretability tests showed a significant difference between the conventional and novel metaphoric senses. For our experiments, we select only the conventional and novel metaphor senses for each word, resulting in 208 sentences (104 conventional and 104 novel metaphors).
GPT4o-metaphors
To test our experiments in a setting that is more controlled (in contrast to VUA) but features more varied sentence lengths, structures and degrees of novelty (in contrast to Lai2009), we construct a synthetic dataset by prompting GPT4o to generate sentences that include conventional and novel metaphoric senses. We do not assume that LLMs are able to generate ideal novel metaphors, or that this dataset is a benchmark of any kind; we are only interested in exploring a new quick setting of potential metaphor novelty annotations. We prompt GPT-4o to generate 5 verbs and 5 nouns that can be used in a metaphoric sense. For each word, we prompt the model again to generate 10 different sentences using the target word in a conventional metaphor sense, and 10 different sentences in a novel metaphor sense. This results in a dataset of 200 sentences, 100 contain conventional metaphors, and 100 contain novel metaphors.
@@ 4.1 Surprisal of a target word
In information theory shannon1948, the information content of an event $x$ with probability $p(x)$ is defined as: $I(x)=-\log p(x)$. In the context of LMs trained to predict the next token in a sequence, this quantity is called surprisal and computed for a $w_{i}$ in a sequence such that: $\mathrm{Surprisal}(w_{i})=-\log p(w_{i}\mid w_{<i})$ [^2] We use $\log$ of base $e$ for all surprisal/perplexity computations in our study..
@@
Naturally, the method of computing word-level surprisal from LMs is a rough approximation process, and design choices of this method can bias the results pimentel-meister-2024-compute; lesci-etal-2025-causal. We identify a couple of choices in the details of this method that can affect  [^3] We find differences in the correlation values when these choices are altered. the final results of our study. Mainly, subword tokenisation is the reason for such methodological confounds.
Firstly, the goal is to get the probability of a target word (e.g. “arrested” in Figure 1); however, most LMs produce probabilities per tokens (subwords) (e.g. [‘‘Ġarrest’’, ‘‘ed’’]) rather than words. And the method of aligning these tokens from the list of sentence tokens to specific words becomes a methodological choice. We choose to align the tokens to words by precomputing the offsets of the exact [^4] We make sure no punctuation or any other characters outside of the word boundary are included inside the offsets. target words in their corresponding sentences, and then searching for the minimal span of tokens in the LM’s tokenisation of the input sentence covering the target word’s offsets. This usually results in a proper alignment (e.g. ‘‘Ġarrested’’ for “arrested”), with very rare cases when the last token is attached to a punctuation (e.g. [‘‘Ġindivid’’, ‘‘ual’’, ‘‘ism,’’] for “individualism”). We also apply the corrections made by pimentel-meister-2024-compute; oh-schuler-2024-leading that address the problem of leading whitespaces in tokenisations of most causal LMs.
@@
We try to address this inherent weakness of word-level surprisal computed from causal models, which does not take the right context of the metaphoric word into account. This is an issue when sentences do not end in the metaphorical word, but feature various positions of the metaphorical word as in naturally occurring corpus data (see Section 3.1), and also in some cases of the GPT-4o-metaphors dataset. To address this, we experiment with a new strategy for computing word-level surprisal autoregressively, while introducing the full sentence context: for each metaphor word, we input a prompt (as illustrated in Figure 1) that indicates a cloze test and show the whole sentence to the model except the metaphor word (replaced by a blank space), and then spell out the whole sentence again as if the model is solving the cloze test. We measure the surprisal of the target word at the second occurrence of the sentence, ensuring that the left and right context are included in the conditional probability of the metaphor word.

### v2

@@ Abstract
Novel metaphor comprehension involves complex semantic processes and linguistic creativity, making it an interesting task for studying language models (LMs). This study investigates whether surprisal, a probabilistic measure of predictability in LMs, correlates with different metaphor novelty datasets. We analyse surprisal from 16 LM variants on corpus-based and synthetic metaphor novelty datasets. We explore a cloze-style surprisal method that conditions on full-sentence context. Results show that LMs yield significant moderate correlations with scores/labels of metaphor novelty. We further identify divergent scaling patterns: on corpus-based data, correlation strength decreases with model size (inverse scaling effect), whereas on synthetic data it increases (Quality–Power Hypothesis). We conclude that while surprisal can partially account for annotations of metaphor novelty, it remains a limited metric of linguistic creativity. [^1] Code and data are publicly available: https://github.com/OmarMomen14/surprisal-metaphor-novelty.
@@
Omar Momen,   Emilie Sitter, Berenike Herrmann and Sina Zarrieß CRC 1646 – Linguistic Creativity in Communication Faculty of Linguistics and Literary Studies Bielefeld University, Germany {omar.hassan,emilie.sitter,berenike.herrmann,sina.zarriess}@uni-bielefeld.de
@@ 1 Introduction
Recent advances in neural language modelling have led to a renewed interest in studying linguistic creativity, an aspect of language that was notoriously challenging for traditional NLP systems. A well-known instance of creative language is metaphor, which arises through mapping of a source domain–a lexical unit’s literal meaning–onto a target domain–its figurative meaning-Lakoff and Johnson (1980). Yet, not all metaphors are equally novel or creative as many of such mappings are highly conventionalized, such as in the famous example “to attack an argument,” where the mapping from the source domain “WAR” onto the target domain “ARGUMENT” corresponds to a conventional sense of the verb “attack” which can be found in dictionaries. Other metaphors, as in “The arrested water” (Figure 1) establish a more unconventional or novel mapping that is considered creative.
@@
Figure 1: A sentence from VUA-ratings with metaphor-novelty ratings (n) and surprisal measures (s) from GPT2-base. The top part shows the direct-surprisal measure, while the bottom part shows the cloze-surprisal measure.
@@
The distinction between conventional and novel metaphors is well established in theory Bowdle and Gentner (2005); Gibbs (2005) and studied in experimental work on cognitive processing Arzouan et al. (2007); Lai et al. (2009). Novel metaphors require greater interpretative effort compared to conventional metaphors, as the unfamiliar mapping requires speakers to construct new connections between domains Philip (2016).
@@
However, separating novel metaphors from conventional ones remains a challenging task for human-based annotations Do Dinh et al. (2018); Reimann and Scheffler (2024). While computational methods on metaphor detection and understanding do not even take this distinction into account in most cases (Maudslay and Teufel, 2022), and some existing studies highlight that novel metaphors are more difficult to detect than conventional ones Neidlein et al. (2020); Tong et al. (2024). In this work, we propose to fill this gap and study questions on LMs processing of novel and conventional metaphors, and investigate LM-based metrics correlations to different setups of metaphor novelty annotations.
@@
We draw on a line of research that has studied the ability of LMs to account for effects of difficulty in human sentence processing and that goes back to surprisal theory (Hale, 2001; Levy, 2008). Surprisal is computed with LMs as the negative log probability of a word in context and has been found to provide a robust predictor for human processing difficulty (e.g., of reading times) (Goodkind and Bicknell, 2018; Shain et al., 2024).
@@
However, recent work draws a mixed picture in terms of which LMs can provide the most robust and cognitively plausible predictors. Oh and Schuler (2023b) observe an inverse scaling effect when testing surprisal estimates from GPT-2 models of different sizes, showing that the surprisal computed with smaller model sizes achieved a better fit with human reading times than larger model sizes. Wilcox et al. (2023), on the other hand, train LMs of small and medium size on a range of languages and find that LM quality generally correlates with its psychometric predictive power.
@@
In this paper, we investigate surprisal on metaphor words computed with LMs as a metric of metaphor novelty. We find significant moderate correlations between LMs’ surprisal and different metaphor novelty annotations. We perform our experiments on four different metaphor novelty datasets coming from corpus-based and synthetic setups, using LMs of different sizes and architectures. Most interestingly, we observe effects supporting the "inverse scale" pattern Oh and Schuler (2023b) on the two corpus-based novelty datasets, and contrary patterns supporting the counter-argument, "Quality-Power" hypothesis Wilcox et al. (2023) on the two synthetic datasets. We also investigate the effect of instruction-tuning on surprisal correlation to metaphor novelty. Moreover, we conduct a deeper analysis of the genre splits, revealing that genre, metaphor density and LM perplexity are potential factors underlying the quality of surprisal as a predictor of metaphor novelty. Finally, we introduce a new method of computing surprisal, cloze-surprisal, to include the right context of the word in its conditional probability. We find that this method can boost the correlation of surprisal with metaphor novelty annotations by a few points. In general, our study establishes a promising direction for studying linguistic creativity with LMs and calls for novel measures and datasets that provide systematic annotations of metaphor novelty across genres.
@@ 2.1 Metaphor Annotations
The most common annotation scheme in corpus-based metaphor studies is the Metaphor Identification Procedure Vrije Universiteit (MIPVU) The Pragglejaz Group (2007); Steen et al. (2010), designed as a reliable step-by-step framework for identifying metaphorically used words. This method was used to construct the VU Amsterdam Metaphor Corpus (VUAMC) Steen et al. (2010), a large-scale, genre-balanced corpus containing 186,673 words sampled from fiction, news, academic writing, and conversations splits in the BNC Baby edition. The VUAMC has become a benchmark for metaphor research, inspiring developments in further English corpora Beigman Klebanov et al. (2018); Mohammad et al. (2016) and more recent multilingual efforts Sanchez-Bayona and Agerri (2022); Egg and Kordoni (2022).
Several studies have extended annotation beyond binary literal/metaphoric labels. Mohler et al. (2016) introduced a four-point metaphoricity scale, considering factors such as vividness and familiarity. Reijnierse et al. (2019) proposed the concept of deliberateness, distinguishing metaphors intended to be recognized as metaphors; novel metaphors are typically deliberate under this framework. Direct novelty annotations were first introduced by Parde and Nielsen (2018), who asked annotators to rate metaphorical word pairs from the VUAMC on a 0–3 novelty scale. Do Dinh et al. (2018) proposed a more comprehensive approach considering all metaphor words in VUAMC, and by aggregating ranked novelty judgments into continuous scores. More recently, Reimann and Scheffler (2024) proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries.
Psycholinguistic studies also produce datasets of metaphors, but these are smaller in size and feature controlled synthetic sentences (fixed words, structures, sentence length, etc.), cf. Cardillo et al. (2010, 2012); Roncero and de Almeida (2014). Some of these also created datasets of metaphors that are classified as conventional or novel metaphors Lai et al. (2009); Ahrens et al. (2024).
@@ 2.2 Predictive Powers of Surprisal
Surprisal estimates from LMs have been shown to capture human processing difficulty across multiple behavioural and neural measures. In self-paced reading and eye-tracking studies, token-level surprisals from causal models significantly predict reading times beyond lexical and syntactic factors Goodkind and Bicknell (2018); Wilcox et al. (2020); Oh and Schuler (2023a). Similar effects have been reported for acceptability judgments, where LM surprisals align with human sensitivity to gradient grammaticality Lau et al. (2017); Meister et al. (2021); Hu et al. (2020); Tjuatja et al. (2025).
@@ 2.3 LM-Based Metrics for Metaphor Novelty
Recent work has explored diverse LM-based approaches—prompting, surprisal, embedding similarity and attention patterns—for the analysis of metaphorical language. Aghazadeh et al. (2022) demonstrate that metaphor-relevant information is encoded in mid-layer embeddings of multilingual pre-trained LMs. Ichien et al. (2024) show that GPT-4 generates interpretations of novel literary metaphors that are favoured by human judges over interpretations by college students, suggesting sensitivity to metaphorical meaning beyond lexical overlap. Pedinotti et al. (2021) tested BERT’s masked token probabilities across conventional, novel, and nonsensical metaphors, showing that novel metaphors tend to receive lower probabilities, but distinctions between novelty and deviancy remained unclear. We extend this study towards more recent LMs and broader comparisons between models, datasets, and genres. Djokic et al. (2021) trained a BERT-based classifier to predict novelty scores jointly with the task of metaphor detection.
@@ 3.1 Corpus‐based Datasets
Do Dinh et al. (2018) collected crowd-sourced ratings of metaphor novelty for VUAMC. In their set-up, annotators were presented four sentences containing an MRW (content words only) and asked to select the best (most novel) and worst (most conventional). Each MRW appeared in six different best-worst scaling comparisons. These annotations were then converted into continuous Best-Worst Scaling scores Kiritchenko and Mohammad (2017) in the range of (-1, +1), with -1 being the most conventional and +1 being the most novel. Additionally, they also convert these scores to binary labels using a threshold of 0.5. This results in labelling 353 metaphors as novel out of 15,155 content metaphoric words in VUA.
@@
Reimann and Scheffler (2024) proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries. They applied this method to VUAMC. In particular, they re-annotated the potentially novel metaphors in VUA according to VUA-ratings and Reijnierse et al. (2019)’s metaphor deliberateness annotations (1,160 potentially novel metaphors in total). Their procedure resulted in labelling only 409 content [^2] We exclude non-content words from this dataset to allow comparison with VUA-ratings, which originally excluded non-content words. metaphoric words as novel out of the 1,160 potentially novel metaphors. We assume the remaining metaphors in VUA to be conventional in our study.
@@ 3.2 Synthetic Datasets
Lai et al. (2009) investigated how our brains handle conventional and novel metaphors differently. To this end, they constructed a set of controlled experimental items featuring metaphors with two degrees of novelty. Two linguists selected 104 words, and constructed 4 sentences for each word, according to the Conceptual Metaphor Theory (CMT) Lakoff and Johnson (1980). For each word, the items include one (i) literal use, (i) conventional metaphor, (iii) novel metaphor and (iv) anomalous use, with the target word as the last word in each sentence. Familiarity and interpretability tests showed a significant difference between the conventional and novel metaphoric senses. For our experiments, we select only the conventional and novel metaphor senses for each word, resulting in 208 sentences (104 conventional and 104 novel metaphors).
GPT-4o-metaphors
To test our experiments in a setting that is more controlled (in contrast to VUA) but features more varied sentence lengths, structures and degrees of novelty (in contrast to Lai2009), we construct a synthetic dataset by prompting GPT-4o to generate sentences that include conventional and novel metaphoric senses. We do not assume that LLMs are able to generate ideal novel metaphors, or that this dataset is a benchmark of any kind; we are only interested in exploring a new quick setting of potential metaphor novelty annotations. We prompt GPT-4o to generate 5 verbs and 5 nouns that can be used in a metaphoric sense. For each word, we prompt the model again to generate 10 different sentences using the target word in a conventional metaphor sense, and 10 different sentences in a novel metaphor sense. This results in a dataset of 200 sentences, 100 contain conventional metaphors, and 100 contain novel metaphors.
@@ 4.1 Surprisal of a target word
In information theory Shannon (1948), the information content of an event $x$ with probability $p(x)$ is defined as: $I(x)=-\log p(x)$.
In the context of LMs trained to predict the next token in a sequence, this quantity is called surprisal and computed for a $w_{i}$ in a sequence such that: $\mathrm{Surprisal}(w_{i})=-\log p(w_{i}\mid w_{<i})$ [^3] We use $\log$ of base $e$ for all surprisal/perplexity computations in our study..
@@
Naturally, the method of computing word-level surprisal from LMs is a rough approximation process, and design choices of this method can bias the results Pimentel and Meister (2024); Lesci et al. (2025). We identify a couple of choices in the details of this method that can affect  [^4] We find differences in the correlation values when these choices are altered. the final results of our study. Mainly, subword tokenisation is the reason for such methodological confounds.
Firstly, the goal is to get the probability of a target word (e.g. “arrested” in Figure 1); however, most LMs produce probabilities per tokens (subwords) (e.g. [‘‘Ġarrest’’, ‘‘ed’’]) rather than words. And the method of aligning these tokens from the list of sentence tokens to specific words becomes a methodological choice. We choose to align the tokens to words by precomputing the offsets of the exact [^5] We make sure no punctuation or any other characters outside of the word boundary are included inside the offsets. target words in their corresponding sentences, and then searching for the minimal span of tokens in the LM’s tokenisation of the input sentence covering the target word’s offsets. This usually results in a proper alignment (e.g. ‘‘Ġarrested’’ for “arrested”), with very rare cases when the last token is attached to a punctuation (e.g. [‘‘Ġindivid’’, ‘‘ual’’, ‘‘ism,’’] for “individualism”). We also apply the corrections made by Pimentel and Meister (2024); Oh and Schuler (2024) that address the problem of leading whitespaces in tokenisations of most causal LMs.
@@
We try to address this inherent weakness of word-level surprisal computed from causal models, which does not take the right context of the metaphoric word into account. This is an issue when sentences do not end in the metaphorical word, but feature various positions of the metaphorical word as in naturally occurring corpus data (see Figure 4), and also in some cases of the GPT-4o-metaphors dataset. To address this, we experiment with a new strategy for computing word-level surprisal autoregressively, while introducing the full sentence context: for each metaphor word, we input a prompt (as illustrated in Figure 1) that indicates a cloze test and show the whole sentence to the model except the metaphor word (replaced by a blank space), and then spell out the whole sentence again as if the model is solving the cloze test. We measure the surprisal of the target word at the second occurrence of the sentence, ensuring that the left and right context are included in the conditional probability of the metaphor word.